# Quantum Dynamics with Stochastic Non-Hermitian Hamiltonians

### v1

@@ Abstract
Note that the trace of the unnormalized Liouvillian acting on any state reads $\mathrm{Tr}(\tilde{\mathcal{L}}[\hat{\rho}])=(4\gamma\Gamma^{2}-2\Gamma)\mathrm{Tr}(\hat{\Pi}\hat{\rho})$. As $\mathrm{Tr}(\hat{\Pi}\hat{\rho})\geq 0$, the dynamics is trace-decreasing when $04\gamma\Gamma^{2}-2\Gamma<$ 0 and trace-increasing when $04\gamma\Gamma^{2}-2\Gamma>$ 0. Therefore, the dynamics exhibits a transition from TD to TI at $\gamma^{*}=\frac{1}{2\Gamma}$. At this critical value, the nonlinear term vanishes: the dynamics is Completely Positive and Trace Preserving (CPTP), and its generator admits the standard GKSL form, with jump operator $\hat{\Pi}$ and rate $\Gamma$. When the Hamiltonian is Hermitian ($\Gamma=0$), the dynamics is trivially unitary. The noise can be used to tune the success rate, i.e. the number of trajectories with no jumps. For TD dynamics, this success rate corresponds to the trace $\mathrm{Tr}(\tilde{\rho}_{t})\propto\exp((4\gamma\Gamma^{2}-2\Gamma)t)$. If the noise increases this decay is slower, thus, a stronger noise $\gamma$ can be used to mitigate the decay of the success rate, especially near the critical value $\gamma^{*}$.
Every state of a qubit can be expressed in the Bloch sphere as $\hat{\rho}_{t}=\frac{1}{2}(\mathbb{1}+\mathbf{r}_{t}\cdot\hat{\boldsymbol{\sigma}})$ Nielsen and Chuang (2010). Expressing the Bloch vector $\mathbf{r}_{t}=(x_{t},y_{t},z_{t})$ in spherical coordinates $(r_{t},\theta_{t},\phi_{t})$, the purity (5) evolves for any state $\hat{\rho}_{t}$ as
@@
Note that this expression shows a change in behavior at the TD-TI transition $\Gamma\gamma=\frac{1}{2}$. When $\Gamma\gamma<\frac{1}{2}$, the purifying states are in the north hemisphere of the Bloch ball, while for $\Gamma\gamma>\frac{1}{2}$, these states are shifted to the southern hemisphere, see Fig. 1.
@@
We now investigate the spectral properties of the SDQ Liouvillian (8). The phase diagrams in Fig. 2 have three distinct phases: at weak noise $\gamma$, the notions of $\mathcal{PT}$ broken $\Gamma/J>2$ and unbroken $\Gamma/J<2$ symmetry of the NH Hamiltonian govern the dynamics, while at large noise strength, we see a transition to the TI dynamics at $\gamma^{*}$. We refer to these phases as $\mathcal{PT}$ broken ($\mathcal{PT}$b), $\mathcal{PT}$ unbroken ($\mathcal{PT}$u) and Noise Induced (NI). Figure 2(a) shows the dissipative gap $\Delta$ as a function of the noise strength $\gamma$ and the decay rate $\Gamma$. In the $\mathcal{PT}$u phase, the dissipative gap is very small, and thus the convergence to the stable steady state is very slow. In the $\mathcal{PT}$b phase the gap is larger, therefore the steady state is reached faster than in the $\mathcal{PT}$u phase. In the NI phase, the gap is very large, ensuring fast convergence to the stable steady state [cf. Eq. (6)]. Note that the gap is always smaller around the transition to the NI phase, which implies that the GKSL dynamics have slower convergence to the stable steady state. Fig. 2(b) shows the maximum imaginary part of the eigenvalues of the Lindbladian, which dictate the oscillatory behavior of the dynamics. Deep in the $\mathcal{PT}$b and NI phases, the imaginary part vanishes, implying that the dynamics is not oscillatory. However, it is nonzero in the $\mathcal{PT}$u phase. This fact, along with the very small dissipative gap, shows what features of the $\mathcal{PT}$u phase survive the application of classical noise.
@@
We now characterize the non-degenerate stable steady state $\hat{\rho}^{\textsc{s}}$. Figure 2(c) shows the $z$ component of the Bloch vector of $\hat{\rho}^{\textsc{s}}$, which reads $z(\hat{\rho}^{\textsc{s}})=-\frac{\lambda_{0}(\lambda_{0}-A)}{4+\lambda_{0}(\lambda_{0}-A)}$ (cf. App. B), where $\lambda_{0}$ denotes the eigenvalue with the largest real part and the constant $A=\frac{\Gamma}{J}(\gamma\Gamma-1)$. In the $\mathcal{PT}$u phase, the $z$ component is small, and the steady state is close to the completely mixed state. Thus, when a very small noise is added to the $\mathcal{PT}$ symmetric NH Hamiltonian, the time evolution leads to a highly mixed state at very long times. In the $\mathcal{PT}$b phase, $z$ is close to unity—so the stable steady state is close to |$\ket{f}$, as losses induced by $\Gamma$ remove all population in the |$\ket{e}$ state. In the NI phase, the steady state is |$\ket{e}$—the state that originally leaked to the ground state. This can be interpreted as a noise-induced transition to stability Horsthemke and Lafever (2006) of the originally unstable state, a feature shared by other noisy dynamics Martinez-Azcona et al. (2023). The $y$ Bloch coordinate phase diagram in Fig. 2(d) shows that the transition from the mixed state to the |$\ket{f}$ (|$\ket{e}$) state happens by acquiring a negative (positive) value of the $y$ coordinate in the Bloch sphere. The analytic expression for this quantity is $y(\hat{\rho}^{\textsc{s}})=\frac{2\lambda_{0}}{4+\lambda_{0}(\lambda_{0}-A)}$ (cf. App. B).
@@
Figure 3: Fidelity between the time-evolved state and the stable steady state as a function of the loss parameter $\Gamma$ and time for noise strength $\gamma J=0.05$. $\mathcal{PT}$u to $\mathcal{PT}$b phase transition (gray dashed) and $\mathcal{PT}$b to NI phase transition (black). Time-scales of the dissipative gap $\Delta^{-1}$ (white dotted) and oscillatory dynamics $2\pi/\omega$ (red dotted).
@@
Let us now turn to the dynamics generated by the master equation (8). The dynamics can be solved numerically in two complementary ways (cf. App. C, D). We characterize the distinguishability between the evolving and stable steady states using the Uhlmann fidelity $F^{\textsc{s}}_{t}=F(\hat{\rho}_{t},\hat{\rho}^{\textsc{s}})$ Uhlmann (1976); Bures (1969), which admits a compact form in a qubit system Jozsa (1994); Hübner (1992) (cf. App. E). Figure 3 shows the evolution of the fidelity as a function of time and the decay parameter $\Gamma$ for a small noise, $\gamma J=0.05$. This value ensures that the three different phases are manifested: (i) for $\Gamma/J<2$, the system is in the $\mathcal{PT}$u phase, and the fidelity to the steady state exhibits the anticipated oscillatory behavior until it vanishes at a time close to the inverse dissipative gap $t\sim\Delta^{-1}$ (white dotted line). The period of these oscillations is perfectly characterized by $2\pi/\omega$ where $\omega=\max_{\nu}\mathrm{Im}(\lambda_{\nu})$, as highlighted by the red dotted line. Note that this oscillatory frequency is affected by the increase of $\Gamma$ as we approach the $\mathcal{PT}$ phase transition, as also observed experimentally in the dissipative qubit Naghiloo et al. (2019). (ii) For $2<\Gamma/J<\frac{1}{2J\gamma}$, the system is in the $\mathcal{PT}$b phase, with a larger dissipative gap [cf. Fig. 2(a)], and thus a faster convergence to the steady state. Interestingly, the oscillatory timescale can be finite outside of the $\mathcal{PT}$u phase as $\omega$ does not vanish identically. In particular, it shows minima at $\Gamma=\frac{1}{3\gamma}$ and $\Gamma=\frac{1}{\gamma}$, as observed in Fig. 2(b). Exactly at the transition $\Gamma=\frac{1}{2\gamma}$, the dynamics is CPTP, and its convergence is slower, as explained by the smaller gap at the transition line. (iii) For $\Gamma>\frac{1}{2\gamma}$, the system is in the NI phase and exhibits the fastest convergence to the steady state. Experimentally, the decay rate is $\Gamma=6.7\;\mu\mathrm{s}^{-1}$ Naghiloo et al. (2019), which upper bounds the noise strength $\gamma<\gamma^{*}\approx 0.074\;\mu\mathrm{s}$ given that no NI phase was observed.
@@
Conclusions. We have considered the time evolution governed by a fluctuating non-Hermitian Hamiltonian describing a quantum mechanical system subject to stochastic gain and loss. The resulting noise-averaged dynamics is described by a novel kind of master equation beyond the GKSL form. We characterized the purity dynamics and found that the stable steady states live in the Liouvillian eigenspace whose eigenvalues have the largest real part. The salient features of such dynamics are manifested in a stochastic dissipative qubit. The addition of noise allows for control over the steady states and the rate of convergence. We find three main dynamical phases: the $\mathcal{PT}$ unbroken and broken regimes, complemented with a noise-induced phase. Our results are amenable to current experimental platforms realizing non-Hermitian evolution. Our findings suggest that engineering fluctuating operators associated with gain and loss may open new avenues for quantum state preparation Verstraete et al. (2009). Our results may also be applied to understand the stability of imaginary time evolution Motta et al. (2020) against classical noise.
@@
Acknowledgements— We thank Oskar Prosniak, Kohei Kawabata, and Martin Plenio for their insightful discussions and Shiue-Yuan Shiau for comments on the manuscript. This research was funded by the Luxembourg National Research Fund (FNR, Attract grant 17132054 and grant 17132054). The work at Los Alamos was supported by the U.S. Department of Energy.
Appendix A Gauge transformations of the nonlinear master equation
@@
We discuss below how trace-increasing dynamics can be changed, such as to conserve or reduce the state trace. For this, we explicitly distinguish the stochastic $\hat{L}_{\textsc{s}}$ and deterministic $\hat{L}_{\textsc{d}}$ parts of the anti-Hermitian part of the Hamiltonian, and add an imaginary deterministic shift $a\in\mathbb{R}$. Namely, we consider
@@
$\partial_{t}\hat{\rho}_{t}=$ $-i[\hat{H}_{0},\hat{\rho}_{t}]-\{\hat{L}_{\textsc{d}}+a\,\hat{\mathbb{1}},\hat{\rho}_{t}\}+\gamma\{\hat{L}_{\textsc{s}},\{\hat{L}_{\textsc{s}},\hat{\rho}_{t}\}\}$
$+2\mathrm{Tr}\big{(}(\hat{L}_{\textsc{d}}+a\,\hat{\mathbb{1}})\hat{\rho}_{t}\big{)}\hat{\rho}_{t}-4\gamma\mathrm{Tr}(\hat{L}_{\textsc{s}}^{2}\hat{\rho}_{t})\hat{\rho}_{t}.$ (11)
@@
$\partial_{t}\hat{\rho}_{t}=$ $-i[\hat{H}_{0},\hat{\rho}_{t}]-\{\hat{L}^{\prime}_{\textsc{d}}-4\gamma b\hat{L}_{\textsc{s}},\hat{\rho}_{t}\}+\gamma\{\hat{L}_{\textsc{s}},\{\hat{L}_{\textsc{s}},\hat{\rho}_{t}\}\}$
$+2\mathrm{Tr}((\hat{L}^{\prime}_{\textsc{d}}-4\gamma b\hat{L}_{\textsc{s}})\hat{\rho}_{t})\hat{\rho}_{t}-4\gamma\mathrm{Tr}(\hat{L}^{2}_{\textsc{s}}\hat{\rho}_{t})\hat{\rho}_{t}.$ (16)
@@
Appendix B Liouvillian spectrum of the Stochastic Dissipative Qubit

### v2

@@ Abstract
The trace of the Liouvillian acting on any state reads $\mathrm{Tr}(\tilde{\mathcal{L}}[\hat{\rho}])=(4\gamma\Gamma^{2}-2\Gamma)\mathrm{Tr}(\hat{\Pi}\hat{\rho})$. Note that it depends on the population of the |$\ket{e}$ state, since $\mathrm{Tr}(\hat{\Pi}\hat{\rho})=\rho_{ee}$. Therefore the dynamics inhibits the state |$\ket{e}$ when $04\gamma\Gamma^{2}-2\Gamma<$ 0 and favors it when $04\gamma\Gamma^{2}-2\Gamma>$ 0. Therefore, the dynamics exhibits a transition at $\gamma^{*}=\frac{1}{2\Gamma}$. At this critical value, the trace of the Liouvillian vanishes, the dynamics is Completely Positive, and Trace Preserving (CPTP), and its generator admits the standard GKSL form, with jump operator $\hat{\Pi}$ and rate $\Gamma$. The dynamics is trivially unitary when the Hamiltonian is Hermitian ($\Gamma=0$). The noise can be used to tune the success rate, i.e., the number of trajectories with no jumps. Assuming that the trace of the unnormalized state gives the success rate, it reads $\mathrm{Tr}(\tilde{\rho}_{t})=e^{-2\Gamma(1-2\gamma\Gamma)t}(c_{0}+JB\int_{0}^{t}e^{-JB\tau}\rho_{ff}(\tau)\mathrm{d}\tau)$ where $JB=2\Gamma(2\gamma\Gamma-1)$. If the noise increases, the overall decay will be slower; thus, a stronger noise $\gamma$ could be used to mitigate the decay of the success rate, especially near the critical value $\gamma^{*}$.”
Every state of a qubit can be expressed in the Bloch sphere as $\hat{\rho}_{t}=\frac{1}{2}(\hat{\mathbb{1}}+\mathbf{r}_{t}\cdot\hat{\boldsymbol{\sigma}})$, where $\hat{\boldsymbol{\sigma}}$ is a vector containing the Pauli matrices Nielsen and Chuang (2010). Expressing the Bloch vector $\mathbf{r}_{t}=(x_{t},y_{t},z_{t})$ in spherical coordinates $(r_{t},\theta_{t},\phi_{t})$, the purity (5) evolves for any state $\hat{\rho}_{t}$ as
@@
Note that this expression shows a change in behavior at the TD-TI transition $\Gamma\gamma=\frac{1}{2}$. When $\Gamma\gamma<\frac{1}{2}$, the purifying states are in the north hemisphere of the Bloch sphere, while for $\Gamma\gamma>\frac{1}{2}$, these states are shifted to the southern hemisphere, see Fig. 1. More details on the steady states are provided in Apps. C, D.
@@
We now investigate the spectral properties of the SDQ Liouvillian (8). The phase diagrams in Fig. 2 have three distinct phases: at weak noise $\gamma$, the notions of $\mathcal{PT}$ broken $\Gamma/J>2$ and unbroken $\Gamma/J<2$ symmetry of the NH Hamiltonian govern the dynamics, while at large noise strength, we see a transition to the TI dynamics at $\gamma^{*}$. We refer to these phases as $\mathcal{PT}$ broken ($\mathcal{PT}$b), $\mathcal{PT}$ unbroken ($\mathcal{PT}$u) and Noise Induced (NI). Figure 2(a) shows the dissipative gap $\Delta$ as a function of the noise strength $\gamma$ and the decay rate $\Gamma$. In the $\mathcal{PT}$u phase, the dissipative gap is very small; thus, the convergence to the stable steady state is very slow. Interestingly, a small residual damping rate was observed experimentally in this phase Naghiloo et al. (2019) (also see App. B). In the $\mathcal{PT}$b phase, the gap is larger. Therefore, the steady state is reached faster than in the $\mathcal{PT}$u phase. In the NI phase, the gap is very large, ensuring fast convergence to the stable steady state [cf. Eq. (6)]. Note that the gap is always smaller around the transition to the NI phase, which implies that the GKSL dynamics has slower convergence to the stable steady state. Fig. 2(b) shows the maximum imaginary part of the eigenvalues of the Lindbladian, which dictate the oscillatory behavior of the dynamics. Deep in the $\mathcal{PT}$b and NI phases, the imaginary part vanishes, implying that the dynamics is not oscillatory. However, it is nonzero in the $\mathcal{PT}$u phase. This fact, along with the very small dissipative gap, shows what features of the $\mathcal{PT}$u phase survive the application of classical noise.
@@
We now characterize the non-degenerate stable steady state $\hat{\rho}^{\textsc{s}}$. Figure 2(c) shows the $z$ component of the Bloch vector of $\hat{\rho}^{\textsc{s}}$, which reads $z(\hat{\rho}^{\textsc{s}})=-\frac{\lambda_{0}(\lambda_{0}-AJ)}{4J^{2}+\lambda_{0}(\lambda_{0}-AJ)}$ (cf. App. C), where $\lambda_{0}$ denotes the eigenvalue with the largest real part and $A=\frac{\Gamma}{J}(\gamma\Gamma-1)$ is a constant. In the $\mathcal{PT}$u phase, the $z$ component is close to zero, and the steady state is close to the maximally mixed state. Thus, when a very small noise is added to the $\mathcal{PT}$ symmetric NH Hamiltonian, the time evolution leads to a highly mixed state at very long times. In the $\mathcal{PT}$b phase, $z$ is close to unity—so the stable steady state is close to |$\ket{f}$, as losses induced by $\Gamma$ remove all population in the |$\ket{e}$ state. In the NI phase, the steady state is |$\ket{e}$—the state that originally leaked to the ground state. This can be interpreted as a noise-induced transition to stability Horsthemke and Lafever (2006) of the originally unstable state, a feature shared by other noisy dynamics Martinez-Azcona et al. (2023). The $y$ Bloch coordinate phase diagram in Fig. 2(d) shows that the transition from the mixed state to the |$\ket{f}$ (|$\ket{e}$) state happens by acquiring a negative (positive) value of the $y$ coordinate in the Bloch sphere. The analytic expression for this quantity is $y(\hat{\rho}^{\textsc{s}})=\frac{2\lambda_{0}J}{4J^{2}+\lambda_{0}(\lambda_{0}-AJ)}$ (cf. App. C).
@@
Figure 3: Fidelity between the time-evolved state and the stable steady state as a function of the decay parameter $\Gamma$ and time for noise strength $\gamma J=0.05$. $\mathcal{PT}$u to $\mathcal{PT}$b phase transition (gray dashed) and $\mathcal{PT}$b to NI phase transition (black). Timescales of the dissipative gap $\Delta^{-1}$ (white dotted) and oscillatory dynamics $2\pi/\omega$ (red dotted).
@@
Let us now turn to the dynamics generated by the master equation (8). The dynamics can be solved numerically in two complementary ways (cf. App. D, E). We characterize the distinguishability between the evolving and stable steady states using the Uhlmann fidelity $F^{\textsc{s}}_{t}=F(\hat{\rho}_{t},\hat{\rho}^{\textsc{s}})$ Uhlmann (1976); Bures (1969), which admits a compact form in a qubit system Jozsa (1994); Hübner (1992) (cf. App. F). Figure 3 shows the evolution of the fidelity as a function of time and the decay parameter $\Gamma$ for a small noise, $\gamma J=0.05$. This value ensures that the three different phases are manifested: (i) for $\frac{\Gamma}{J}<2$, the system is in the $\mathcal{PT}$u phase, and the fidelity to the steady state exhibits the anticipated oscillatory behavior until it vanishes at a time close to the inverse dissipative gap $t\sim\Delta^{-1}$ (white dotted line). The period of these oscillations is perfectly characterized by $2\pi/\omega$ where $\omega=\max_{\nu}\mathrm{Im}(\lambda_{\nu})$, as highlighted by the red dotted line. Note that this oscillatory frequency is affected by the increase of $\Gamma$ as we approach the $\mathcal{PT}$ phase transition, as also observed experimentally in the dissipative qubit Naghiloo et al. (2019). (ii) For $2<\frac{\Gamma}{J}<\frac{1}{2J\gamma}$, the system is in the $\mathcal{PT}$b phase, with a larger dissipative gap [cf. Fig. 2(a)], and thus a faster convergence to the steady state. Interestingly, the oscillatory timescale can be finite outside the $\mathcal{PT}$u phase as $\omega$ does not vanish identically. In particular, it shows minima at $\Gamma=\frac{1}{3\gamma}$ and $\Gamma=\frac{1}{\gamma}$, as observed in Fig. 2(b). Exactly at the transition $\Gamma=\frac{1}{2\gamma}$, the dynamics is CPTP, and its convergence is slower, as explained by the smaller gap at the transition line. (iii) For $\Gamma>\frac{1}{2\gamma}$, the system is in the NI phase and exhibits the fastest convergence to the steady state. Experimentally, the decay rate is $\Gamma=6.7\;\mu\mathrm{s}^{-1}$ Naghiloo et al. (2019), which upper bounds the noise strength $\gamma<\gamma^{*}\approx 0.074\;\mu\mathrm{s}$ given that no NI phase was observed.
@@
Conclusions— We have considered the time evolution governed by a fluctuating non-Hermitian Hamiltonian describing a quantum mechanical system subject to stochastic gain and loss. The resulting noise-averaged dynamics is described by a novel anti-dephasing master equation beyond the GKSL form. We characterized the purity dynamics and found that the stable steady states live in the Liouvillian eigenspace whose eigenvalues have the largest real part. The salient features of such dynamics are manifested in a stochastic dissipative qubit. The addition of noise allows for control over the steady states and the convergence rate. We find three main dynamical phases: the $\mathcal{PT}$ unbroken and broken regimes, complemented with a noise-induced phase. Our results are amenable to current experimental platforms realizing non-Hermitian evolution and provide a framework to capture the effect of noise, as exemplified by the residual damping rate Naghiloo et al. (2019). Our findings suggest that engineering fluctuating operators associated with gain and loss may open new avenues for quantum state preparation Verstraete et al. (2009). Our results may also be applied to understand the stability of imaginary time evolution Motta et al. (2020) against classical noise.
@@
Acknowledgements— We thank Oskar Prośniak, Kohei Kawabata, Martin Plenio, Raphaël Chétrite, and Yogesh Joglekar for their insightful discussions and Juan Garrahan and Shiue-Yuan Shiau for comments on the manuscript. This research was funded by the Luxembourg National Research Fund (FNR, Attract grant QOMPET 15382998 and grant 17132054). The work at Los Alamos was supported by the U.S. Department of Energy.
Appendix A Gauge transformations of the anti-dephasing master equation
@@
Below, we discuss how trace-increasing dynamics can be changed to conserve or reduce the state trace. For this, we explicitly distinguish the stochastic $\hat{L}_{\textsc{s}}$ and deterministic $\hat{L}_{\textsc{d}}$ parts of the anti-Hermitian part of the Hamiltonian, and add an imaginary deterministic shift $a\in\mathbb{R}$. Namely, we consider
@@
$\partial_{t}\hat{\rho}_{t}=$ $-i[\hat{H}_{0},\hat{\rho}_{t}]-\{\hat{L}_{\textsc{d}}+a\,\hat{\mathbb{1}},\hat{\rho}_{t}\}+\gamma\{\hat{L}_{\textsc{s}},\{\hat{L}_{\textsc{s}},\hat{\rho}_{t}\}\}+2\mathrm{Tr}\big{(}(\hat{L}_{\textsc{d}}+a\,\hat{\mathbb{1}})\hat{\rho}_{t}\big{)}\hat{\rho}_{t}-4\gamma\mathrm{Tr}(\hat{L}_{\textsc{s}}^{2}\hat{\rho}_{t})\hat{\rho}_{t}.$ (11)
@@
$\partial_{t}\hat{\rho}_{t}=$ $-i[\hat{H}_{0},\hat{\rho}_{t}]-\{\hat{L}^{\prime}_{\textsc{d}}-4\gamma b\hat{L}_{\textsc{s}},\hat{\rho}_{t}\}+\gamma\{\hat{L}_{\textsc{s}},\{\hat{L}_{\textsc{s}},\hat{\rho}_{t}\}\}+2\mathrm{Tr}((\hat{L}^{\prime}_{\textsc{d}}-4\gamma b\hat{L}_{\textsc{s}})\hat{\rho}_{t})\hat{\rho}_{t}-4\gamma\mathrm{Tr}(\hat{L}^{2}_{\textsc{s}}\hat{\rho}_{t})\hat{\rho}_{t}.$ (16)
@@
Appendix B Experimental Realization of Stochastic Non-Hermitian Hamiltonians
In this Section, we discuss possible experimental platforms for realizing stochastic non-Hermitian Hamiltonians and analyze their advantages and disadvantages. We begin by reviewing the experimental realization of the non-Hermitian qubit in superconducting circuits Naghiloo et al. (2019). In particular, we focus on how noise affects this platform. Then, to show that noise in the anti-Hermitian part of the Hamiltonian does not need to be postulated ad hoc but can be obtained from first principles, we review the formalism by Carmichael Carmichael (1991) in homodyne detection of the degenerate parametric oscillator. Lastly, since tuning the strength of the noise in both of these examples seems quite challenging, we review the setup of the trapped ion non-Hermitian qubit Quinn et al. (2023), which may allow for a tunable noise.
B.1 Noise in the superconducting non-Hermitian qubit
The superconducting non-Hermitian qubit studied in Naghiloo et al. (2019) is built from the three lowest levels of a transmon circuit, which is embedded in a cavity with an impedance mismatch element that allows tuning its density of states, thus modifying the decay of the transmon. In addition, an external magnetic flux is threaded through the SQUID loop of the transmon to tune the decay parameter $\Gamma$ of the effective non-Hermitian system. A Josephson parametric amplifier is used for state tomography and post-selection of quantum jumps.
The ideal (noise-less) behavior of the non-Hermitian qubit is such that, in the $\mathcal{PT}$ unbroken phase, its Hamiltonian’s eigenenergies are purely real and thus describe a purely oscillatory behavior. In contrast, in the $\mathcal{PT}$-broken phase, the eigenvalues become purely imaginary, describing the system states as decaying with a real exponential. However, the experiments show Naghiloo et al. (2019), in the $\mathcal{PT}$ unbroken phase, a residual small damping term $\Gamma_{\mathrm{R}}$ describing a damped oscillation, which is larger than expected from the decay channel of the |$\ket{f}$ state. The authors interpreted this residual decay rate to be associated with the “charge and flux noiseNaghiloo et al. (2019). Our model brings some theoretical understanding of the flux noise.
As explained before, the magnetic flux tunes the non-Hermitian decay rate $\Gamma$, such that noise in the flux corresponds to noise in the anti-Hermitian part. Therefore, our SDQ models flux noise acting on the NH qubit. This noise is unavoidable in the experimental setup. Unfortunately, this also implies that tuning noise in this setup would be difficult. Still, our results show that the $\mathcal{PT}$ unbroken phase in Fig. 2(a) has a small but nonzero, dissipative gap. So, our formalism can serve as a minimal model to capture the effect of noise on the experimental setup.
B.2 The degenerate parametric oscillator
Carmichael derived a stochastic non-Hermitian Hamiltonian for a model of homodyne detection of the degenerate parametric oscillator Carmichael (1991). We briefly review some of its most important features, along with some extensions, for completeness. With this, we aim to show that the stochasticity, particularly noise in the anti-Hermitian part of the Hamiltonian, does not need to arise from uncontrolled interactions with an environment, but can come from a formal treatment of the measurement process.
The physical setup consists of two optical cavities $A$ and $B$ with respective annihilation operators $\hat{a},\;\hat{b}$. The cavity $\hat{a}$ with frequency $\omega$ is prepared in the vacuum state |$\ket{0}$. Carmichael’s setup Carmichael (1991) contains a nonlinear crystal, parametrized by $\lambda$, and radiates squeezed light. However, this parameter is unnecessary to understand the main idea so that we will set it to zero $\lambda=0$. Cavity $\hat{b}$ is prepared in a coherent state |$\ket{\beta}$ and radiates the local oscillator field. Both cavities are leaky and thus modeled by a quantum optical Lindblad master equation with rates $2\kappa$ for the pumped cavity and $2\gamma$ for the local oscillator. The two output fields go through a beam splitter, with a reflection coefficient $R$, before arriving at the detector. The evolution is then split into two different steps: a non-unitary evolution corresponding to no quantum jumps at the detector and a collapse when a photoelectron is emitted.
The model can be simplified to get a pure state evolution of the cavity A, conditioned on the measurement output, given by |$\ket{\psi^{(a)}_{c}(t)}$. To do so, the reflectivity and the local oscillator decay rate are sent to zero ($R\rightarrow 0,\;\gamma\rightarrow 0$) while the local oscillator amplitude is taken to infinity $\beta\rightarrow\infty$, to keep the local oscillator photon flux constant, $f=2\gamma R|\beta|^{2}$. Under these conditions, the unnormalized conditioned state |$\ket{\tilde{\psi}_{c}^{(a)}}$ evolves with the non-Hermitian Hamiltonian
$\hat{H}=\hbar(\omega-i\kappa)\hat{a}^{\dagger}\hat{a}-i\hbar\sqrt{2\kappa f}e^{-i\theta}e^{i\omega t}\hat{a},$ (18)
and is interrupted by collapses
|$\ket{\tilde{\psi}_{c}^{(a)}}\rightarrow\hat{C}\ket{\tilde{\psi}_{c}^{(a)}}=\left(\sqrt{f}e^{i(\theta-\omega t)}+\sqrt{2\kappa}\hat{a}\right)\ket{\tilde{\psi}_{c}^{(a)}},$ (19)
where $\theta$ is the local oscillator phase. The collapses happen at probability $p_{c}(t)~{}=~{}\delta t{\braket{\tilde{\psi}_{c}^{(a)}(t)}{\hat{C}^{\dagger}\hat{C}}{\tilde{\psi}_{c}^{(a)}(t)}}/{\braket{\tilde{\psi}_{c}^{(a)}(t)}{\tilde{\psi}_{c}^{(a)}(t)}}$. Taking the limit $f/(2\kappa)\rightarrow\infty$, the wavefunction undergoes an infinite number of infinitesimal collapses in any time interval, such that the jump process reduces to white noise. After some algebra, taking the time between successive collapses as $\tau_{n+1}\sim f^{-1}$ and the total number of collapses $m$ in the increment of time $\delta t$ from a Gaussian distribution Carmichael (1991), we find that the conditioned, unnormalized state of the source |$\ket{\tilde{\psi}_{c}^{(a)}(t)}$ evolves with a Schrödinger equation
$i\hbar\partial_{t}\ket{\tilde{\psi}_{c}^{(a)}(t)}=\hat{H}_{\xi}(t)\ket{\tilde{\psi}_{c}^{(a)}(t)},$ (20)
where $\hat{H}_{\xi}(t)$ is a stochastic, non-Hermitian Hamiltonian given by
$\hat{H}_{\xi}(t)$ $=\hbar(\omega-i\kappa)\hat{a}^{\dagger}\hat{a}+i\hbar\left(\sqrt{2\kappa}\braket{\psi_{c}^{(a)}(t)}{\hat{X}_{\theta}}{\psi_{c}^{(a)}(t)}+\xi_{t}\right)e^{-i\theta}\sqrt{2\kappa}\hat{a},$ (21)
where we introduced the operator $\hat{X}_{\theta}=e^{i\theta}\hat{a}^{\dagger}+e^{-i\theta}\hat{a}$ and $\xi_{t}$ is a Gaussian white noise. Note that Carmichael Carmichael (1991) uses the interaction picture with respect to $\omega\hat{a}^{\dagger}\hat{a}$, leading to a purely anti-Hermitian Hamiltonian. Here, for clarity, we do not use the interaction picture and thus obtain a Hamiltonian with both a Hermitian and an anti-Hermitian part. Interestingly, a similar stochastic non-Hermitian Hamiltonian has been proposed by Piñol et al. Piñol et al. (2024) for a two-level system instead of an oscillator, i.e. with $\hat{\sigma}_{\pm}$ instead of $\hat{a},\hat{a}^{\dagger}$.
B.3 Trapped Ion realization
Recently, a non-Hermitian trapped-ion qubit has been realized Quinn et al. (2023). The experimental realization uses a single ${}^{40}\mathrm{Ca}^{+}$ ion in a linear-Paul trap Sherman et al. (2013). The two energy levels used to build a qubit are |=|$\ket{\uparrow}=\ket{m=+5/2}$ and |=|$\ket{\downarrow}=\ket{m=+3/2}$ within the meta-stable $D_{5/2}$ manifold. The hopping between the states $J\sigma_{x}$ uses resonant radio frequency pulses at the qubit frequency. To engineer the decay, the state $\ket{\downarrow}$ is coupled to the short-lived $P_{3/2}$ state |$\ket{A}$, which decays primarily to the $S_{1/2}$ ground state with rate $\gamma_{g}$. The coupling to the |$\ket{A}$ state is achieved with $\pi$-polarized light of pulse strength $J_{A}$. For $\gamma_{g}\gg J_{A}$, the auxiliary level |$\ket{A}$ may be adiabatically eliminated Brion et al. (2007); Muga et al. (2008) and through post-selection, an effective Non-Hermitian Qubit is obtained for the levels {$\{\ket{\downarrow},\ket{\uparrow}\}$. The effective decay rate of $\ket{\downarrow}$ is then $\gamma=J_{A}^{2}/\gamma_{g}\ll\gamma_{g}$. The pulse strength may then be modulated in time with some external and tunable white noise $J_{A}(1+\sqrt{2\gamma^{\prime}}\xi_{t}^{\prime})$, to achieve a fluctuating decay rate.
A limitation of the experimental setup is that $5.87\%$ of the population of the $P_{3/2}$ state decays back to the $D_{5/2}$ manifold Quinn et al. (2023), limiting the time that the effective non-Hermitian description is valid over one or two Rabi oscillations. This might still be a suitable platform to observe our predicted results. Indeed, one of the main results of the Stochastic Dissipative Qubit is the presence of a noise-induced phase. Figure 3 of the main text shows that the dissipative gap time scale in the NI phase is much shorter than the Rabi oscillation time, which implies that the NI phase should be experimentally observable in this trapped-ion platform.
Appendix C Liouvillian spectrum of the Stochastic Dissipative Qubit